OpenStudy (anonymous):

y varies directly as the square of x. When x=3, y= 54. Find y when x=8.

OpenStudy (anonymous):

y=kx Variation equation 54=3k k=18 So plugging that in y=(18)(8) y=144

OpenStudy (anonymous):

thank you

OpenStudy (kropot72):

\[y=kx ^{2}\]

OpenStudy (anonymous):

My bad, @kropot72 got me beat

OpenStudy (kropot72):

\[y=kx ^{2}\] \[54=k(3)^{2}\] \[k=\frac{54}{3^{2}}=6\] When x = 8 \[y=6\times (8)^{2}\]

OpenStudy (anonymous):

\[54=k(3^2)\] \[6=k\] \[y=6*(8^2)\] 384